Solve: 5|1-4x|-15=0
Answer by jim_thompson5910(35256)

5|1-4x|-15=0

5|1-4x| = 15

|1-4x| = 15/5

|1-4x| = 3

1-4x = 3 or 1-4x = -3

-4x = 3-1 or -4x = -3-1

-4x = 2 or -4x = -4

x = 2/(-4) or x = -4/(-4)

x = -1/2 or x = 1